# FitCF: A Framework for Automatic Feature Importance-guided Counterfactual Example Generation

### v1

@@ Abstract
Counterfactual examples are widely used in natural language processing (NLP) as valuable data to improve models, and in explainable artificial intelligence (XAI) to understand model behavior. The automated generation of counterfactual examples remains a challenging task even for large language models (LLMs), despite their impressive performance on many tasks. In this paper, we first introduce ZeroCF, a faithful approach for leveraging important words derived from feature attribution methods to generate counterfactual examples in a zero-shot setting. Second, we present a new framework, FitCF [^1] Our code will be publicly available once the paper is published., which further verifies aforementioned counterfactuals by label flip verification and then inserts them as demonstrations for few-shot prompting, outperforming two state-of-the-art baselines. Through ablation studies, we identify the importance of each of FitCF’s core components in improving the quality of counterfactuals, as assessed through flip rate, perplexity, and similarity measures. Furthermore, we show the effectiveness of LIME and Integrated Gradients as backbone attribution methods for FitCF and find that the number of demonstrations has the largest effect on performance. Finally, we reveal a strong correlation between the faithfulness of feature attribution scores and the quality of generated counterfactuals.
@@
Qianli Wang1,2   Nils Feldhus1,2,6   Simon Ostermann1,3,5 Luis Felipe Villa-Arenas1,2,4   Sebastian Möller1,2   Vera Schmitt1,2 1German Research Center for Artificial Intelligence (DFKI) 2Technische Universität Berlin   3Saarland Informatics Campus   4Deutsche Telekom 5 Centre for European Research in Trusted AI (CERTAIN) 6BIFOLD – Berlin Institute for the Foundations of Learning and Data {firstname.lastname}@dfki.de
@@ 1 Introduction
The advent of increasingly complex and opaque LLMs has triggered a critical need for explainability and interpretability of such models. Counterfactuals, which are minimally edited inputs that yield different predictions compared to reference inputs Miller (2019); Ross et al. (2021); Madsen et al. (2022) are widely used in XAI and NLP. Applications include creating new data points for improving models in terms of performance Kaushik et al. (2020); Sachdeva et al. (2024) and robustness Gardner et al. (2020); Ross et al. (2021) and understanding the black-box nature of models Wu et al. (2021); Wang et al. (2024). Crowd-sourcing counterfactuals can be costly, inefficient, and impractical Chen et al. (2023), particularly in specialized domain such as medicine. LLM-based counterfactual generation offers a more efficient and scalable alternative. Despite advancements in counterfactual generation techniques and the demonstrated versatility of LLMs across tasks Wu et al. (2021); Bhan et al. (2023); Li et al. (2024), the efficacy of LLMs in producing high-quality counterfactuals in a zero-shot setting, as well as the effective construction of valid counterfactuals as demonstrations to enable few-shot prompting, remains an open question Bhattacharjee et al. (2024b). Additionally, the combination of widely used interpretability methods, with the goal to exploit their combined benefits, has been insufficiently explored within XAI research Treviso et al. (2023); Baeumel et al. (2023); Bhan et al. (2023).
@@
Secondly, we evaluate ZeroCF and FitCF on two NLP tasks - news topic classification and sentiment analysis - using two baselines, Polyjuice Wu et al. (2021) and FIZLE Bhattacharjee et al. (2024b). The automatic evaluation employs three automated metrics: Label flip rate, fluency, and edit distance. Both ZeroCF and FitCF significantly outperform Polyjuice, with ZeroCF surpassing FIZLE in most cases and FitCF consistently exceeding both baselines and ZeroCF.
@@ 2 Related Work
MICE generates contrastive edits that change the prediction to a given contrast prediction Ross et al. (2021). Polyjuice uses a fine-tuned GPT-2 Radford et al. (2019) to specify the type of edit needed to generate counterfactual examples Wu et al. (2021). DISCO Chen et al. (2023) uses the GPT-3 fill-in-the-blank mode Brown et al. (2020), which is not available in most open-source LLMs Chen et al. (2023). Bhattacharjee et al. (2024a) identify the latent features in the input text and the input features associated with the latent features to generate counterfactual examples, which is criticized due to the additional level of complexity with no significant performance gain Delaunay et al. (2024). FIZLE Bhattacharjee et al. (2024b) shares the most similarity with FitCF and uses LLMs as pseudo-oracles to generate counterfactuals with the assistance of LLM-generated important words in a zero-shot setting.
@@
Recent works have explored the possibility to combine different XAI methods. Wang et al. (2021) propose a feature importance-aware attack, which disrupts important features that consistently influence the model’s decisions. Gressel et al. (2023) identify perturbations in the feature space to produce evasion attacks. Treviso et al. (2023) present the framework, CREST, to generate counterfactual examples by combining rationalization with span-level masked language modeling. Krishna et al. (2023) employ various post-hoc explanations for rationalization, extending beyond counterfactuals, in contrast to CREST. Bhan et al. (2023) propose a method to determine impactful input tokens with respect to generated counterfactual examples. In contrast, FitCF uses feature importance to guide counterfactual example generation.
@@ 3.1 ZeroCF
Bhattacharjee et al. (2024b) introduced FIZLE, which generates counterfactuals in a zero-shot setting by prompting the LLM with important words identified by the LLM itself. However, these extracted words may be unfaithful or hallucinated Li et al. (2023) [^2] Applying Llama3-8B with FIZLE on AG News, we find that for 64.5% of the instances, a subset of generated important words is hallucinated, i.e., absent from the original input.. To address this limitation, we propose ZeroCF (Figure 2; examples are provided in Table 7), which relies on the most attributed words based on feature attribution scores determined by various explanation methods for the predictions of a BERT model fine-tuned on the target dataset. Feature importance involves determining how significant an input feature is for a given output Madsen et al. (2022), which we find to enhance the counterfactual generation process (§6.1).
@@ 3.2 FitCF
In order to diversify demonstration selection An et al. (2023); Zhang et al. (2023) and construct demonstrations automatically, we first convert all instances from the dataset $\mathcal{D}$ into sentence embeddings using SBERT [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, and then apply $k$-means clustering on these sentence embeddings to form $c$ clusters [^6] Clustering visualizations are given in Appendix C.. Afterwards, we select a total of $k$ instances which are closest to the centroid of each cluster [^7] Selected examples and their corresponding counterfactuals for a given instance are provided in Appendix D.. In such a way, we diversify the demonstrations, potentially mitigating any misleading effects caused by ZeroCF, which may produce flawed counterfactuals. Finally, ZeroCF is employed to generate counterfactuals for the $k$ selected instances using simple heuristics.
@@
Subsequently, in order to validate the generated counterfactuals and to prevent incorrect counterfactuals from misleading the LLM Turpin et al. (2023), we employ the same BERT model $\mathcal{M_{D}}$3.1) to make predictions on $k$ generated counterfactuals $\mathcal{C}=\{\tilde{x}_{1},\tilde{x}_{2},...,\tilde{x}_{k}\}$ and the original input $\mathcal{X}=\{x_{1},x_{2},...,x_{k}\}$ individually and assess whether the labels are inconsistent:
$\forall i\in\{1,2,\cdots,k\}:\hat{y}_{x_{i}}=\mathcal{M_{D}}(x_{i})$ (4)
$\forall i\in\{1,2,\cdots,k\}:\hat{y}_{\tilde{x}_{i}}=\mathcal{M_{D}}(\tilde{x}_{i})$ (5)
The generated counterfactuals $\tilde{x}_{i}$, where the predicted labels remain consistent $\hat{y}_{\tilde{x}_{i}}=\hat{y}_{x_{i}}$, are excluded from the demonstrations for further process to ensure the validity of the generated counterfactuals. In the end, we obtain $m$ counterfactuals, where $m\leq k$. To maintain a consistent number of demonstrations ($\ell$) for each input, if $m<\ell$, additional examples are iteratively selected based on their proximity to the cluster centroid, until the required number of demonstrations is achieved.
@@ 3.3 Considerations for Choice of Models
In ZeroCF, feature attributions are generated for a BERT model’s predictions, based on which important words are then extracted(§3.1). Moreover, in FitCF, the same BERT model serves as a label flip verifier (§3.2). We emphasize that any model capable of performing classification tasks effectively can be used as a label flip verifier or for generating feature attribution scores [^8] For encoder-only architectures like the BERT model employed in our study, tools like Ferret Attanasio et al. (2023) can be used to derive feature attribution scores (§4.4). For encoder-decoder or decoder-only architectures, tools like Inseq Sarti et al. (2023) can generate such scores..
@@ 4.1 Baselines
We employ the following two approaches as baselines for FitCF.
@@
Polyjuice allows users to control perturbation types and deploys a GPT-2999Although Polyjuice utilizes a relatively small model, GPT-2, for generating counterfactuals, we fairly consider it a suitable baseline for FitCF, since the deployed GPT-2 is fine-tuned on a counterfactual example dataset. to generate counterfactuals by framing the task as a conditional text generation problem Wu et al. (2021).
@@
FIZLE employs an LLM to identify important words and prompts the LLM with these words in a zero-shot setting to generate counterfactuals Bhattacharjee et al. (2024b).
@@ 4.2 Dataset
Following Nguyen et al. (2024); Bhattacharjee et al. (2024b), we demonstrate the validity of ZeroCF and FitCF by applying them to two NLP tasks: News topic classification and sentiment analysis [^10] Details on label distributions and example instances from the datasets used can be found in Appendix E..
@@ 5.1 Automatic Evaluation
The Soft Label Flip Rate (SLFR) measures the frequency at which newly perturbed examples alter the original label to a different label Ge et al. (2021); Nguyen et al. (2024); Bhattacharjee et al. (2024a). For a dataset with $N$ instances, we calculate SLFR as follows:
@@
Following Wang et al. (2023); Nguyen et al. (2024); Bhattacharjee et al. (2024b), we deploy GPT-2 to calculate PPL in our experiments due to its proven effectiveness in capturing such text distributions.
@@
$TS=\frac{1}{N}\sum_{i=1}^{N}\frac{d(x_{i},\tilde{x}_{i})}{|x_{i}|}$ (6)
@@ 5.2.2 Effect of Number of Demonstrations
In FitCF, as $c$ clusters are obtained through clustering, and due to the difficulty and complexity of counterfactual example generation, we set the number of demonstrations to twice the number of clusters ($2c$) for each dataset (§3.2; Figure 4), which results in 10 demonstrations for AG News and 8 for SST2, respectively. To examine the effect of the number of demonstrations and assess the necessity of doubling the number of demonstrations to $2c$, we further evaluate the quality of counterfactual examples generated by FitCF, with the number of demonstrations set to the number of clusters ($c$).

### v3

@@ Abstract
Counterfactual examples are widely used in natural language processing (NLP) as valuable data to improve models, and in explainable artificial intelligence (XAI) to understand model behavior. The automated generation of counterfactual examples remains a challenging task even for large language models (LLMs), despite their impressive performance on many tasks. In this paper, we first introduce ZeroCF, a faithful approach for leveraging important words derived from feature attribution methods to generate counterfactual examples in a zero-shot setting. Second, we present a new framework, FitCF [^1] Code is available at: https://github.com/qiaw99/FitCF. FitCF stands for “Feature Importance-guided Counterfactual Example Generation”., which further verifies aforementioned counterfactuals by label flip verification and then inserts them as demonstrations for few-shot prompting, outperforming three state-of-the-art baselines. Through ablation studies, we identify the importance of each of FitCF’s core components in improving the quality of counterfactuals, as assessed through flip rate, perplexity, and similarity measures. Furthermore, we show the effectiveness of LIME and Integrated Gradients as backbone attribution methods for FitCF and find that the number of demonstrations has the largest effect on performance. Finally, we reveal a strong correlation between the faithfulness of feature attribution scores and the quality of generated counterfactuals, which we hope will serve as an important finding for future research in this direction.
@@
Qianli Wang1,2,22footnotemark: 2   Nils Feldhus1,2,6   Simon Ostermann2,4,5 Luis Felipe Villa-Arenas1,2,3   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin 2German Research Center for Artificial Intelligence (DFKI)  3Deutsche Telekom 4Saarland Informatics Campus  5Centre for European Research in Trusted AI (CERTAIN) 6BIFOLD – Berlin Institute for the Foundations of Learning and Data 22footnotemark: 2 Correspondence: qianli.wang@tu-berlin.de
@@ 1 Introduction
The advent of increasingly complex and opaque LLMs has triggered a critical need for explainability and interpretability of such models. Counterfactuals, which are minimally edited inputs that yield different predictions compared to reference inputs Miller (2019); Ross et al. (2021); Madsen et al. (2022) are widely used in XAI and NLP. Applications include creating new data points for improving models in terms of performance Kaushik et al. (2020); Sachdeva et al. (2024) and robustness Gardner et al. (2020); Ross et al. (2021), and understanding the black-box nature of models Wu et al. (2021); Wang et al. (2024a, b). Crowd-sourcing counterfactuals can be costly, inefficient, and impractical Chen et al. (2023), particularly in a specialized domain such as medicine. LLM-based counterfactual generation offers a more efficient and scalable alternative. Despite advancements in counterfactual generation techniques and the demonstrated versatility of LLMs across tasks Wu et al. (2021); Bhan et al. (2023a); Li et al. (2024), the efficacy of LLMs in producing high-quality counterfactuals in a zero-shot setting, as well as the effective construction of valid counterfactuals as demonstrations to enable few-shot prompting, remains an open question Bhattacharjee et al. (2024a). Additionally, the combination of widely used interpretability methods, with the goal to exploit their combined benefits, has been insufficiently explored within XAI research Treviso et al. (2023); Baeumel et al. (2023); Bhan et al. (2023a).
@@
Secondly, we evaluate ZeroCF and FitCF on two NLP tasks - news topic classification and sentiment analysis - using three baselines, Polyjuice Wu et al. (2021), BAE Garg and Ramakrishnan (2020) and FIZLE Bhattacharjee et al. (2024a). The automatic evaluation employs three automated metrics: Label flip rate, fluency, and edit distance. Both ZeroCF and FitCF significantly outperform Polyjuice and BAE, with ZeroCF surpassing FIZLE in most cases and FitCF consistently exceeding three state-of-the-art baselines and ZeroCF.
@@ 2 Related Work
MICE generates contrastive edits that change the prediction to a given contrast prediction Ross et al. (2021). Polyjuice uses a fine-tuned GPT-2 Radford et al. (2019) to specify the type of edit needed to generate counterfactual examples Wu et al. (2021). DISCO Chen et al. (2023) uses the GPT-3 fill-in-the-blank mode, which is not available in most open-source LLMs Chen et al. (2023). Tigtec Bhan et al. (2023b) utilizes local feature importance to identify words that significantly influence a model’s prediction and masks these important words using a fine-tuned masking model. Bhattacharjee et al. (2024b) identify the latent features in the input text and the input features associated with the latent features to generate counterfactual examples, which is criticized due to the additional level of complexity with no significant performance gain Delaunay et al. (2024). FIZLE Bhattacharjee et al. (2024a) shares the most similarity with FitCF and uses LLMs as pseudo-oracles to generate counterfactuals with the assistance of LLM-generated important words in a zero-shot setting.
@@
Recent works have explored the possibility to combine different XAI methods. Wang et al. (2021) propose a feature importance-aware attack, which disrupts important features that consistently influence the model’s decisions. Gressel et al. (2023) identify perturbations in the feature space to produce evasion attacks. Treviso et al. (2023) present the framework, CREST, to generate counterfactual examples by combining rationalization with span-level masked language modeling. Krishna et al. (2023) employ various post-hoc explanations for rationalization, extending beyond counterfactuals, in contrast to CREST. Bhan et al. (2023a) propose a method to determine impactful input tokens with respect to generated counterfactual examples. In contrast, FitCF uses feature importance to guide counterfactual example generation.
@@ 3.1 ZeroCF
Bhattacharjee et al. (2024a) introduced FIZLE, which generates counterfactuals in a zero-shot setting by prompting the LLM with important words identified by the LLM itself. However, these extracted words may be unfaithful or hallucinated Li et al. (2023) [^2] Applying Llama3-8B with FIZLE on AG News, we find that for 64.5% of the instances, a subset of generated important words is hallucinated, i.e., absent from the original input.. To address this limitation, we propose ZeroCF (Figure 2; examples are provided in Table 7), which relies on the most attributed words based on feature attribution scores determined by various explanation methods for the predictions of a BERT model fine-tuned on the target dataset. Feature importance involves determining how significant an input feature is for a given output Madsen et al. (2022), which we find to enhance the counterfactual generation process (§6.1).
@@ 3.2 FitCF
In order to diversify demonstration selection An et al. (2023); Zhang et al. (2023) and construct demonstrations automatically, we first convert all instances from the dataset $\mathcal{D}$ into sentence embeddings using SBERT [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, and then apply $k$-means clustering on these sentence embeddings to form $k$ clusters [^6] Clustering visualizations are given in Appendix C., where $k$ does not necessarily correspond to the exact number of predefined dataset labels. Afterwards, we select a total of $c$ instances which are closest to the centroid of each cluster [^7] Selected examples and their corresponding counterfactuals for a given instance are provided in Appendix D.. In such a way, we diversify the demonstrations, potentially mitigating any misleading effects caused by ZeroCF, which may produce flawed counterfactuals. Finally, ZeroCF is employed to generate counterfactuals for the $c$ selected instances using simple heuristics.
@@
Subsequently, in order to validate the generated counterfactuals and to prevent incorrect counterfactuals from misleading the LLM Turpin et al. (2023), we employ the same BERT model $\mathcal{M_{D}}$3.1) to make predictions on $c$ generated counterfactuals $\mathcal{C}=\{\tilde{x}_{1},\tilde{x}_{2},...,\tilde{x}_{c}\}$ and the original input $\mathcal{X}=\{x_{1},x_{2},...,x_{c}\}$ individually and assess whether the labels are inconsistent:
$\forall i\in\{1,2,\cdots,c\}:\hat{y}_{x_{i}}=\mathcal{M_{D}}(x_{i})$ (4)
$\forall i\in\{1,2,\cdots,c\}:\hat{y}_{\tilde{x}_{i}}=\mathcal{M_{D}}(\tilde{x}_{i})$ (5)
The generated counterfactuals $\tilde{x}_{i}$, where the predicted labels remain consistent $\hat{y}_{\tilde{x}_{i}}=\hat{y}_{x_{i}}$, are excluded from the demonstrations for further process to ensure the validity of the generated counterfactuals. In the end, we obtain $m$ counterfactuals, where $m\leq c$. To maintain a consistent number of demonstrations ($\ell$) for each input, if $m<\ell$, additional examples are iteratively selected based on their proximity to the cluster centroid, until the required number of demonstrations is achieved. Moreover, given we select $\ell$ demonstrations with $\ell>>k$, the influence of the number of clusters, $k$, on the final performance is diminished.
@@ 3.3 Considerations for Choice of Models
In ZeroCF, feature attributions are generated for a BERT model’s predictions, based on which important words are then extracted(§3.1). Moreover, in FitCF, the same BERT model serves as a label flip verifier (§3.2). BERT is chosen as our design choice, because it is a representative encoder language model with a strong efficiency-performance balance. We emphasize that any model capable of performing classification tasks effectively can be used as a label flip verifier or for generating feature attribution scores. For encoder-only architectures like the BERT model employed in our study, tools like Ferret Attanasio et al. (2023) can be used to derive feature attribution scores (§4.4). For encoder-decoder or decoder-only architectures, tools like Inseq Sarti et al. (2023) can generate such scores.
@@ 4.1 Baselines
We employ three approaches as baselines for ZeroCF and FitCF, including FIZLE that achieves state-of-the-art performance in zero-shot counterfactual generation, which we aim to enhance.
BAE
BAE is an adversarial attack method that employs BERT to perturb input text by replacing masked words Garg and Ramakrishnan (2020).
@@
Polyjuice allows users to control perturbation types and deploys a GPT-2888Although Polyjuice utilizes a relatively small model, GPT-2, for generating counterfactuals, and BAE uses BERT to replace words based on embeddings, we consider both of them suitable baseline methods for FitCF. This is because the deployed GPT-2 is fine-tuned on a counterfactual example dataset. Furthermore, FIZLE relies on zero-shot prompting and achieves state-of-the-art performance. to generate counterfactuals by framing the task as a conditional text generation problem Wu et al. (2021).
@@
FIZLE employs an LLM to identify important words and prompts the LLM with these words in a zero-shot setting to generate counterfactuals Bhattacharjee et al. (2024a).
@@ 4.2 Dataset
Following Nguyen et al. (2024); Bhattacharjee et al. (2024a), we demonstrate the validity of ZeroCF and FitCF by applying them to two NLP tasks: News topic classification and sentiment analysis [^9] Details on label distributions and example instances from the datasets used can be found in Appendix E..
@@ 5.1 Automatic Evaluation
The Soft Label Flip Rate (SLFR) measures the frequency at which newly perturbed examples alter the original label to a different label Ge et al. (2021); Nguyen et al. (2024); Bhattacharjee et al. (2024b). For a dataset with $N$ instances, we calculate SLFR as follows:
@@
Following Wang et al. (2023); Nguyen et al. (2024); Bhattacharjee et al. (2024a), we deploy GPT-2 to calculate PPL in our experiments due to its proven effectiveness in capturing such text distributions.
@@
$TS=\frac{1}{N}\sum_{i=1}^{N}\frac{d(x_{i},\tilde{x}_{i})}{|x_{i}|}$
@@ 5.2.1 Effect of Important Words
Approach Dataset AG News ($PPL=95.72$) SST2 ($PPL=309.53$)
Model Method SLFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ SLFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
Polyjuice GPT2 - 18.60% 121.76 0.50 29.00% 258.32 0.71
BAE BERT - 19.50% 168.44 0.12 47.00% 367.06 0.09
FIZLE Llama3-8B - 93.50% 123.67 0.61 95.50% 202.22 0.52
Qwen2.5-32B - 49.00% 53.07 1.14 86.80% 167.51 0.66
Qwen2.5-72B - 21.50% 84.09 0.22 92.00% 257.91 0.43
ZeroCF (Ours) Llama3-8B Gradient 93.50% 102.56 0.38 97.50% 239.15 0.46
IG 95.50% 109.09 0.27 99.50% 222.51 0.42
LIME 97.50% 107.72 0.39 97.00% 264.91 0.42
SHAP 98.00% 99.08 0.27 94.00% 204.76 0.46
Qwen2.5-32B Gradient 68.00% 62.63 2.10 70.50% 205.06 0.48
IG 51.00% 60.45 0.76 91.00% 222.57 0.64
LIME 56.00% 63.75 0.84 90.50% 576.59 0.62
SHAP 55.50% 61.68 0.79 93.00% 191.00 0.60
Qwen2.5-72B Gradient 16.67% 74.19 0.21 88.50% 263.47 0.34
IG 24.50% 92.47 0.22 92.00% 281.10 0.46
LIME 23.00% 72.73 0.71 85.00% 289.20 0.30
SHAP 25.00% 73.92 0.74 86.50% 319.60 0.22
FitCF (Ours) Llama3-8B Gradient 94.50% 86.90 0.21 99.80% 159.57 0.47
IG 96.00% 87.67 0.23 100.00% 161.88 0.48
LIME 95.50% 75.15 0.19 100.00% 151.22 0.48
SHAP 94.00% 260.57 0.21 100.00% 157.36 0.49
Qwen2.5-32B Gradient 56.00% 62.97 0.73 89.00% 214.25 0.51
IG 57.50% 57.01 0.68 90.50% 221.64 0.49
LIME 56.00% 57.45 0.79 89.50% 174.34 0.52
SHAP 62.00% 57.64 0.78 89.50% 157.09 0.52
Qwen2.5-72B Gradient 77.00% 62.13 0.99 96.00% 595.71 0.38
IG 42.00% 63.54 0.33 95.00% 207.55 0.39
LIME 45.00% 61.54 0.35 96.50% 240.94 0.41
SHAP 38.96% 67.28 0.34 96.50% 590.94 0.39
Table 1: Automatic evaluation results of counterfactuals generated by Polyjuice, BAE, FIZLE, ZeroCF, and FitCF with Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B using Soft Label Flip Rate (SLFR), Perplexity (PPL), and Textual Similarity (TS) on AG News and SST2. Bold faced values indicate for each approach, which feature importance method is the best performing according to the respective metric.
@@ 5.2.2 Effect of Number of Demonstrations
In FitCF, as $k$ clusters are obtained through clustering, and due to the difficulty and complexity of counterfactual example generation, we set the number of demonstrations to twice the number of clusters for each dataset ($\ell=2k$; §3.2), which results in 10 demonstrations for AG News and 8 for SST2, respectively (Figure 4). To examine the effect of the number of demonstrations and assess the necessity of doubling the number of demonstrations to $2k$, we further evaluate the quality of counterfactual examples generated by FitCF, with the number of demonstrations set to the number of clusters ($\ell=k$).